# Faster Motion Planning via Restarts

### v1

@@ 4.4 The wide search
The wide search starts being effective as soon as the running time of the first algorithm exceeds a threshold $\nabla$ that already provide a significant probability that the original algorithm succeeds, as testified by the following lemma.
@@
Let $T_{j}=\nabla\left\lceil{j/\beta}\right\rceil$. If we run the first algorithm for $T_{j}$ time, then the first $\left\lceil{j/\beta}\right\rceil$ algorithms in the wide search are going each to be run at least $T_{j}/\left\lceil{j/\beta}\right\rceil=\nabla$ time. The probability that all these algorithms fail to successfully terminate is at most $p_{j}=(1-\beta)^{\left\lceil{j/\beta}\right\rceil}\leq\exp(-j)$. The total running time of the simulation until this point in time is bounded by
@@ 4.5 Caching runs
Another natural approach that utilizes pause/resume operations, is, when running any simulation, suspending running algorithms when their TTL is met instead of terminating then, thus conserving some of the effort expended by running them. Now, whenever the algorithm needs to be run for a certain time threshold $t_{i}$, we first check if there is a suspended run that can be resumed to achieve the desired threshold. To avoid encumbering the system by flooding it with suspended runs (as the wide search does), we limit the number of runs stored in the cache, and kill the “lighter” runs, that is, runs whose runtime is lower, when that number is exceeded.
This recycling trades some of randomness of original scheme for computational efficiency, as long runs of alg are not started from scratch.
@@ 5 Experiments
Here, we present the experiments conducted to evaluate the advantages and weaknesses of the different restarting strategies when applied to SBMP algorithms. We start by recounting the compared methods and providing important details about the experimental setup, followed by the description of the experiments. For every experiment we provide the base details pertaining to the environment, robot, SBMP algorithm, and motion planning task, the results, and a short discussion.
Our objective is to highlight the difference in performance caused by using restarting strategies. Since this framework is especially suitable for parallel implementations of SBMP algorithms, providing a simple and powerful way to harness any number of computation nodes, we compare the different strategies against a parallel-OR benchmark, described in Section 2. To demonstrate the framework’s applicability to the single process model, we first show results on using a single thread before proceeding to the main body of experiments.
Our experiments are all done in simulation, and include a $6$DOF fixed-base manipulator, a $6$DOF free-flying object, and mobile planar robots. The first two experiments are conducted in 3D environments that give rise to RRT queries with a high-variance runtime distribution. The second experiment is conducted in a 2D maze that give rise to PRM queries with a low-variance runtime distribution. These experiments aim to show the results of using stochastic resetting in scenarios with different distributions.
The last experiment, only constituting a proof-of-concept, is conducted in a 2D environment specifically designed to have two distinct paths. The goal of this experiment is to show that, in some cases, restarting strategies are expected to find shorter paths than those found by the parallel-OR strategy.
@@ 5.1.1 Implementation details
Our restarting simulator was implemented in C++ in Linux using the standard system process mechanism. In order to not overwhelm the system with too many child processes running simultaneously. Thus, as a precaution, if the system we used has a CPU with $t$ threads (we had $t=16$ and $t=128$ in our experiments), our simulation used $(3/4)t$. We used the SBMP algorithm implementations from the open source Parasol Planning Library (PPL) [Par25]. For both RRT and PRM, default PPL parameter values were used.
Every single execution of a strategy, i.e., finding a solution for a given motion planning problem, possibly by running many copies of some algorithm alg, was capped at $3000$ seconds. This failure threshold was selected to allow us to run a significant number of experiments on hard instances (relative to the hardware used). Throughout the tables summarizing the runtime results of the experiments the number of failed applications of the strategy are highlighted red. Also, the descriptive statistics in all tables refer only to successful runs.
@@ 5.1.3 List of strategies
    Parallel: The parallel-OR model. Run $\mathcalb{p}$ independent copies of alg using $\mathcalb{p}$ threads (in parallel) until one of them succeeds. See Section 2.2 for more details on the parallel-OR model.
@@
    Wide: A shared counter can be queried for a copy of alg that needs to be advanced, which could either be a paused process or a new copy. Each of the $\mathcalb{p}$ threads queries the counter, and advances the process by a number of seconds equivalent to its current total runtime. New copies of alg are run for one second. If the process does not successfully terminates during this time it is paused, and the thread queries the counter again. See Section 4.4 for more details on the wide search strategy.
@@
    Random $\zeta(2)$: Every thread draws a random number $t$ from the $\zeta(2)$ distribution, and runs a copy of alg with TTL $t$. In practice this means running $\mathcalb{p}$ independent copies of random $\zeta(2)$ search. See Section 4.3.1 for more details on the random $\zeta(2)$ distribution strategy.
@@
    Counter+Cache: Same as the counter strategy, except that the shared counter responds to queries with a TTL $t$, and possibly a paused process $P$ that can be resumed in order to achieve a run of alg with the target TTL. The querying thread then either starts a new run of alg with TTL $t$, or resumes $P$ until it wither succeeds or reaches $t$ seconds of runtime. If the TTL is reached, the current process is pushed to the cache, a fixed-size priority queue based on the runtime. The cache size is $\mathcalb{P}$, where $\mathcalb{P}$ is the maximum number of threads on the current architecture (so 16 on L12, and 128 on L96). See Section 4.5 for more details on the X + cache strategy.
Note that we do not compare all strategies in every experiment. The subset of strategies used depends on the tested hypothesis. Also, to keep the number of experiments we perform under control, we only use caching in tandem with the counter search strategy, although our experimentation suggest it might be helpful in many cases.
@@ 5.2 Inputs
    Simple Passage (Figure 5.1). We use simple passage environment as a benchmark for high-variance cases, as both the geometries involved and the task are intuitive.
    • Environment: A $10\times 10\times 10$ bounding box split by a wall with a single square passage with side-length 1.7 in its center.
    • Robot: $6$DOF or free flying $2\times 1\times 1$ rectangular prism.
@@
    • Environment: A table with a fixed manipulator at its center, and a double shelf is in front of the manipulator. The top shelf contains 3 objects, a cylindrical object, and two rectangular prisms.
    • Robot: $6$DOF manipulator.
    • Task: Move from a neutral position outside the shelf to a configuration with the robot’s end-effector inside the top shelf and it’s arm between two objects.
@@
    Figure 5.2: The Manipulator input. An illustration of the experiment described in Section 5.3.2. The start and goal configurations are colored red and green respectively.
@@
    • Environment: 3D hollow oblong shape with a single thin funnel shaped entrance with the thin part of the funnel inside the trap.
    • Robot: $6$DOF free-flying elongated cylinder.
@@
    Figure 5.3: An illustration of the experiment described in Section 5.3.3. The start and goal configurations are colored blue and red respectively.
@@
    Figure 5.4: An illustration of the experiment described in Section 5.4. The start and goal configurations are colored blue and red respectively.
@@
    • Environment: A 2D bounding box of size $100\times 20$ with a wall crossing it from left to right. The wall has two passages, a narrow passage close to the left end, and a wider passage between the wall and the right end of the bounding box.
    • Robot: $3$DOF rectangular robot
@@
    We used this input to demonstrate that our strategies leads to shorter solutions, see Table 5.5.
@@ 5.3.1 Basic benchmark
The first experiment, run on input Simple Passage, serves as a comparison benchmark for the rest of the section. We run all of the strategies described above, as well as a single thread RRT for comparison. The simplicity of this input is intended to demonstrate the strategies without other considerations getting in the way, such as the difficulty of the input itself, or unexpected behaviors of the motion planner that rise over more complicated inputs. To that end we run this experiment on L12.
This experiment includes a simple narrow passage environment and a basic $6$DOF free-flying object, see Simple Passage, which are enough to create a high-variance instance of the motion planning problem, as a “lucky” sample with the right orientation may result in an almost immediate solution, but, if some configurations with a “bad” orientation are sampled next to the passage’s opening, catastrophic runtimes may be encountered.
@@
Results and discussion. The results, summarized in Table 5.1, show the super-linear improvement of using additional threads in high-variance motion planning problems. The high variance phenomena is also apparent in the high failure rate ($33\%$) of the single thread method (i.e., the runtime exceeded 3000 seconds in a single try).
The remarkable advantage of the different restarting strategies over the parallel-OR model is also evident, with the slowest restarting strategy (wide search) giving an average runtime almost $5$ times faster than that of the parallel strategy. The large variance and gap between the min and the max in Table 5.1 are further evidence of the high variance.
@@ 5.3.2 Fixed base manipulator
In this experiment, run on the Manipulator input, a manipulator resembling a UR5 tasked with reaching into a cluttered shelf, serves two purposes. On top of showcasing the advantage of the restarting strategies over the parallel-OR model, it is the only experiment where we demonstrate the effectiveness of the restarting strategies on a “plain” single process architecture, and the difference between fixed-TTL strategies to ours. Recall that the optimal strategy is always a fixed-TTL strategy. See Observation 3.3.
Intuitively, this motion planning problem, described bellow in greater detail, contains a narrow c-space passage as the manipulator must find it’s way between the objects on the shelf.
@@
Results and discussion. The results, summarized in Table 5.2, show the superiority of the different restarting strategies in all scenarios. In Table 5.2 (a), we see that a simple run of RRT trying to solve the shelf motion planning task has a success rate of between $6\%-36\%$ that of the restarting strategies. The difference between the restarting strategies is also remarkable, with the worst performing strategy, random $\zeta(2)$ achieving a success rate only $\sim 42\%$ that of the top performer counter + cache. The descriptive statistics other than success rate are obviously meaningless in this case.
In Table 5.2 (b) and (c) we see that as the number of thread increases, The differences between the non-fixed strategies shrink. The mean and median runtimes are relatively similar on both L12 and L96 machines, and no single non-fixed strategy shows an unambiguous advantage over the others. The difference between the parallel-OR model and these strategies remains quite large. On the L12 machine we stopped the experiments with the parallel strategy to save time, after it became clear that it will not produce meaningful statistical results due to low success rates, and on the L96 machine it maintains RRT’s high variance runtime distribution, and achieves substantially slower runtimes on average.
In Table 5.2 (c) we see that fixed-TTL strategies perform better on this problem as the threshold approaches 3, surpassing the non-fixed strategies. This conforms with the theoretical analyses and calculations we performed by following Lemma 3.4 and using an empirical distribution to approximate the ground truth $\mathcal{D}$ of the runtime of RRT on this instance.
@@ 5.3.3 Mobile robot
Results and discussion. In Table 5.3 summarizing the results of this experiment we, once more, see the clear advantage of all restarting strategies over the parallel-OR model, whose runs on L12 we, again, stopped after it became clear that the low success rate will render the statistical results insignificant. Wide search stands out in this experiment with relatively poor results. We hypothesize that due to the difficulty of the problem (for RRT), the number of paused processes quickly became too large for the machine, even L96, to efficiently handle, thus leading to poor results. In particular, the hardware of L96 seems to be unfriendly to the multitude of processes that the wide search creates, and seems to under-perform for this strategy. This phenomena is also evident in our other experiments.
@@ 5.4 Low variance experiment
This is the only experiment in which we use PRM, as it is more suitable for this problem, and enabled us to run a full set of experiments much quicker. The purpose of this experiment is to emphasize that the restarting strategies advantage stems directly from the high variance of the problem, completing our presentation of the general framework.
@@
Results and discussion. The results, summarized in Table 5.4, show that, as expected, the parallel-OR method outperforms all of the restarting strategies. This fits the theoretical analysis, as the parallel strategy does not waste time on short runs of PRM, and its runs reach more advanced stages of PRM’s run in which success probability is meaningful.
@@ 5.5 Path length experiment
In this experiment we measure and report only the lengths of the paths returned by the algorithm. All of the numbers in Table 5.5 refer only to that quantity.
@@
6 Conclusions
In this paper, we have demonstrated the applicability of the stochastic resetting framework to SBMP algorithms. We have provided a comprehensive introduction to the theory and intuition behind it, and experimentally verified its effectiveness in motion planning instances that give rise to high variance runtime distributions. We have also presented several novel restarting strategies that are asymptotically optimal and easy to parallelize.
How to use. We provide an open source implementation of the framework [AH25a]. We emphasize that converting an existing program $P$ to use the framework is quite easy. One needs to create a simple wrapper script that:
@@ 6 Conclusions
Thus, using our framework can be implemented in less than 20 minutes. In particular, it might be useful as a tool for harnessing parallelism in a multi-core machine, in a situation where the existing implementation is single thread.
When to use. Generally speaking the speedup techniques shines when the descriptive complexity of the solution is “small” – the optimal solution is a simple motion involving few parts. Conversely, the speedup technique fails miserably if the motion is complicated (such as in the Maze input, see Table 5.4).
Future research. A natural direction is to try and break complicated motion planning problems into short chunks where our techniques help, and combine the solutions together to get a global complicated solution faster.

### v2

@@ 3.4 The wide search with pause/resume operations
The wide search starts being effective as soon as the running time of the first algorithm exceeds a threshold $\nabla$ that already provides a significant probability that the original algorithm succeeds, as testified by the following lemma.
@@
Let $T_{j}=\nabla\left\lceil{j/\beta}\right\rceil$. If we run the first algorithm for $T_{j}$ time, then the first $\left\lceil{j/\beta}\right\rceil$ algorithms in the wide search are going to be run at least $T_{j}/\left\lceil{j/\beta}\right\rceil=\nabla$ time. The probability that all these algorithms fail to successfully terminate is at most $p_{j}=(1-\beta)^{\left\lceil{j/\beta}\right\rceil}\leq\exp(-j)$. The total running time of the simulation until this point in time is bounded by
@@ 3.5 Caching runs
Another natural approach that utilizes pause/resume operations is, when running any simulation, suspending running algorithms when their TTL is met instead of terminating them, thus conserving some of the effort expended by running them. Now, whenever the algorithm needs to be run for a certain time threshold $t_{i}$, we first check if there is a suspended run that can be resumed to achieve the desired threshold. To avoid encumbering the system by flooding it with suspended runs (as the wide search does), we limit the number of runs stored in the cache, and kill the “lighter” runs, that is, runs whose runtime is lower, when that number is exceeded.
This recycling trades some of the randomness of the original scheme for computational efficiency, as long runs of alg are not started from scratch.
@@ 4 Experiments
Here, we present the experiments conducted to evaluate the advantages and weaknesses of the different restarting strategies when applied to SBMP algorithms. We start by recounting the compared methods and providing important details about the experimental setup, followed by the description of the experiments. For every experiment, we provide the base details about the environment, robot, SBMP algorithm, and motion planning task, the results, and a short discussion.
Our objective is to highlight the difference in performance caused by using restarting strategies. Since this framework is especially suitable for parallel implementations of SBMP algorithms, providing a simple and powerful way to harness any number of computation nodes, we compare the different strategies against a parallel-OR benchmark, described in Section 1.2. To demonstrate the framework’s applicability to the single process model, we first show results on using a single thread before proceeding to the main body of experiments.
Our experiments are all done in simulation, and include a 6 degrees of freedom (DOF) fixed-base manipulator, a  [^6] DOF free-flying object, and mobile planar robots. The first two experiments are conducted in 3D environments that give rise to RRT queries with a high-variance runtime distribution. The second experiment is conducted in a 2D maze that gives rise to PRM queries with a low-variance runtime distribution. These experiments aim to show the results of using stochastic resetting in scenarios with different distributions.
The last experiment, only constituting a proof-of-concept, was conducted in a 2D environment specifically designed to have two distinct paths. The goal of this experiment is to show that, in some cases, restarting strategies are expected to find shorter paths than those found by the parallel-OR strategy.
@@ 4.1.1 Implementation details
[…]
Every single execution of a strategy, i.e., finding a solution for a given motion planning problem, possibly by running many copies of some algorithm alg, was capped at 300030003000 seconds. This failure threshold was selected to allow us to run a significant number of experiments on hard instances (relative to the hardware used). Throughout the tables summarizing the runtime results of the experiments, the number of failed applications of the strategy is highlighted in red. Also, the descriptive statistics in all tables refer only to successful runs.
@@ 4.1.3 List of strategies
    Parallel: The parallel-OR model. Run $\mathcalb{p}$ independent copies of alg using $\mathcalb{p}$ threads (in parallel) until one of them succeeds. See Section 1.2.2 for more details on the parallel-OR model.
@@
    Wide: A shared counter can be queried for a copy of alg that needs to be advanced, which could either be a paused process or a new copy. Each of the $\mathcalb{p}$ threads queries the counter, and advances the process by a number of seconds equivalent to its current total runtime. New copies of alg are run for one second. If the process does not successfully terminate during this time, it is paused, and the thread queries the counter again. See Section 3.4 for more details on the wide search strategy.
@@
    Random $\zeta(2)$: Every thread draws a random number $t$ from the $\zeta(2)$ distribution, and runs a copy of alg with TTL $t$. In practice, this means running $\mathcalb{p}$ independent copies of a random $\zeta(2)$ search. See Section 3.3.1 for more details on the random $\zeta(2)$ distribution strategy.
@@
    Counter+Cache: Same as the counter strategy, except that the shared counter responds to queries with a TTL $t$, and possibly a paused process $P$ that can be resumed to achieve a run of alg with the target TTL. The querying thread then either starts a new run of alg with TTL $t$, or resumes $P$ until it either succeeds or reaches $t$ seconds of runtime. If the TTL is reached, the current process is pushed to the cache, a fixed-size priority queue based on the runtime. The cache size is $\mathcalb{P}$, where $\mathcalb{P}$ is the maximum number of threads on the current architecture (so 16 on L12, and 128 on L96). See Section 3.5 for more details on the X + cache strategy.
Note that we do not compare all strategies in every experiment. The subset of strategies used depends on the tested hypothesis. Also, to keep the number of experiments we perform under control, we only use caching in tandem with the counter search strategy, although our experimentation suggests it might be helpful in many cases.
@@ 4.2 Inputs
    Simple Passage (Figure 4.1). We use a simple passage environment as a benchmark for high-variance cases, as both the geometries involved and the task are intuitive.
    • Environment: A $\times 10\times 10$ 10 bounding box split by a wall with a single square passage with side-length 1.7 in its center.
    • Robot:  [^6] DOF or free flying 2×1×12$\times$ 1$\times$ 12 × 1 × 1 rectangular prism.
@@
    • Environment: A table with a fixed manipulator at its center, and a double shelf is in front of the manipulator. The top shelf contains 3 objects: a cylindrical object and two rectangular prisms.
    • Robot:  [^6] DOF manipulator.
    • Task: Move from a neutral position outside the shelf to a configuration with the robot’s end-effector inside the top shelf and its arm between two objects.
@@
    Figure 4.2: The Manipulator input. An illustration of the experiment described in Section 4.3.2. The start and goal configurations are colored red and green, respectively.
@@
    • Environment: 3D hollow oblong shape with a single thin funnel-shaped entrance with the thin part of the funnel inside the trap.
    • Robot:  [^6] DOF free-flying elongated cylinder.
@@
    Figure 4.3: An illustration of the experiment described in Section 4.3.3. The start and goal configurations are colored blue and red, respectively.
@@
    Figure 4.4: An illustration of the experiment described in Section 4.4. The start and goal configurations are colored blue and red, respectively.
@@
    • Environment: A 2D bounding box of size 100×20100$\times$ 20100 × 20 with a wall crossing it from left to right. The wall has two passages, a narrow passage close to the left end, and a wider passage between the wall and the right end of the bounding box.
    • Robot:  [^3] DOF rectangular robot
@@
    We used this input to demonstrate that our strategies lead to shorter solutions, see Table 4.5.
@@ 4.3.1 Basic benchmark
The first experiment, run on input Simple Passage, serves as a comparison benchmark for the rest of the section. We run all of the strategies described above, as well as a single thread RRT for comparison. The simplicity of this input is intended to demonstrate the strategies without other considerations getting in the way, such as the difficulty of the input itself or unexpected behaviors of the motion planner that arise over more complicated inputs. To that end, we run this experiment on L12.
This experiment includes a simple narrow passage environment and a basic  [^6] DOF free-flying object, see Simple Passage, which are enough to create a high-variance instance of the motion planning problem, as a “lucky” sample with the right orientation may result in an almost immediate solution, but, if some configurations with a “bad” orientation are sampled next to the passage’s opening, catastrophic runtimes may be encountered.
@@
Results and discussion. The results, summarized in Table 4.1, show the super-linear improvement of using additional threads in high-variance motion planning problems. The high variance phenomenon is also apparent in the high failure rate ($\%3$) of the single thread method (i.e., the runtime exceeded 3000 seconds in a single try).
The remarkable advantage of the different restarting strategies over the parallel-OR model is also evident, with the slowest restarting strategy (wide search) giving an average runtime almost 555 times faster than that of the parallel strategy. The large variance and gap between the min and the max in Table 4.1 are further evidence of the high variance.
@@ 4.3.2 Fixed base manipulator
In this experiment, run on the Manipulator input, a manipulator resembling a UR5 tasked with reaching into a cluttered shelf serves two purposes. On top of showcasing the advantage of the restarting strategies over the parallel-OR model, it is the only experiment where we demonstrate the effectiveness of the restarting strategies on a “plain” single process architecture, and the difference between fixed-TTL strategies to ours. Recall that the optimal strategy is always a fixed-TTL strategy. See Observation 2.3.
Intuitively, this motion planning problem, described below in greater detail, contains a bottleneck in the form of a small set of configurations that must be sampled or found for a valid motion to be generated (known as a narrow passage) as the manipulator must find its way between the objects on the shelf.
@@
[…]
In Table 4.2 (b) and (c), we see that as the number of threads increases, the differences between the non-fixed strategies shrink. The mean and median runtimes are relatively similar on both L12 and L96 machines, and no single non-fixed strategy shows an unambiguous advantage over the others. The difference between the parallel-OR model and these strategies remains quite large. On the L12 machine, we stopped the experiments with the parallel strategy to save time, after it became clear that it would not produce meaningful statistical results due to low success rates, and on the L96 machine, it maintains RRT’s high variance runtime distribution, and achieves substantially slower runtimes on average.
In Table 4.2 (c), we see that fixed-TTL strategies perform better on this problem as the threshold approaches 3, surpassing the non-fixed strategies. This conforms with the theoretical analyses and calculations we performed by following Lemma 2.4 and using an empirical distribution to approximate the ground truth $\mathcal{D}$ of the runtime of RRT on this instance.
@@ 4.3.3 Mobile robot
Results and discussion. Table 4.3, summarizing the results of this experiment, shows again the clear advantage of all restarting strategies over the parallel-OR model, whose runs on L12 we, again, stopped after it became clear that the low success rate would render the statistical results insignificant. Wide search stands out in this experiment, with relatively poor results. We hypothesize that due to the difficulty of the problem (for RRT), the number of paused processes quickly became too large for the machine, even L96, to efficiently handle, thus leading to poor results. In particular, the hardware of L96 seems to be unfriendly to the multitude of processes that the wide search creates, and seems to underperform for this strategy. This phenomenon is also evident in our other experiments.
@@ 4.4 Low variance experiment
This is the only experiment in which we use PRM, as it is more suitable for this problem, and enabled us to run a full set of experiments much quicker. The purpose of this experiment is to emphasize that the restarting strategies’ advantage stems directly from the high variance of the problem, completing our presentation of the general framework.
@@
Results and discussion. The results, summarized in Table 4.4, show that, as expected, the parallel-OR method outperforms all of the restarting strategies. This fits the theoretical analysis, as the parallel strategy does not waste time on short runs of PRM, and its runs reach more advanced stages of PRM’s run in which the success probability is meaningful.
@@ 4.5 Path length experiment
In this experiment, we measure and report only the lengths of the paths returned by the algorithm. All of the numbers in Table 4.5 refer only to that quantity.
@@
5 How to use our implementation (for SBMP)
Stochastic restarting strategies are practical and impactful tools that can enhance the performance of SBMP algorithms. Below, we provide recommendations for applying these methods.
How to use. We provide an open source implementation of the framework [AH25]. We emphasize that converting an existing program $P$ to use the framework is quite easy. One needs to create a simple wrapper script that:
@@ 5 How to use our implementation (for SBMP)
Thus, using our framework can be implemented in less than 20 minutes. In particular, it might be useful as a tool for harnessing parallelism in a multi-core machine, in a situation where the existing implementation is single-threaded.
When to use. Generally speaking, the speedup techniques shine when the descriptive complexity of the solution is “small” – the optimal solution is a simple motion involving few parts. Conversely, the speedup technique fails miserably if the motion is complicated (such as in the Maze input, see Table 4.4). If using $N$ threads, using $N-1$ with restarting strategies and the remaining thread with “plain” RRT avoids catastrophic runtimes, while taking into account the possibility of a low-variance instance.
6 Conclusions
In this paper, we have demonstrated the applicability of the stochastic resetting framework to SBMP algorithms. We have provided a comprehensive introduction to the theory and intuition behind it, and experimentally verified its effectiveness in motion planning instances that give rise to high-variance runtime distributions. We have also presented several novel restarting strategies that are asymptotically optimal and easy to parallelize.
Future research. A natural direction is to try and break complicated motion planning problems into short chunks where our techniques help, and combine the solutions to get a global complicated solution faster.